Suppose for a commodity X, demand and supply curves are given by the equations

                             (1)         Qd = 5/P

                             (2)         Qs = -4+P


  What is the equilibrium price?

 What is the equilibrium quantity?


1
Expert's answer
2021-07-23T05:22:21-0400

Solution:

At equilibrium: Quantity demanded (Qd) = Quantity supplied (Qs)

Qd = Qs

"\frac{5}{P} = -4 +P"

Multiply both sides by P:

5 = -4P + P2

-4P + P2 = 5

-4P + P2 – 5 = 0

P2 – 4P – 5 = 0

P (P – 4) – 5 = 0

Therefore, P = 5


Equilibrium price = 5


Substitute in either the quantity demanded equation or quantity supplied equation to derive the equilibrium quantity:

Qd = 5/P

Qd = 5/5 = 1

Qs = -4 + P

Qs = -4 + 5 = 1

Equilibrium quantity = 1 unit